# Multivariate Probabilistic Time Series Forecasting with Correlated Errors

### v3

@@ A.3 Training Procedure
Compute used All models in the paper were trained in an Anaconda environment with access to one AMD Ryzen Threadripper PRO 5955WX CPU and four NVIDIA RTX A5000 GPU (24 GB of memory).
Batch size We adopt the approach of GPVar [2] by using $B=20$ time series in a sample slice and a batch size of 16. Given that our data sampler selects one slice of time series as a batch instead of sampling 16 slices simultaneously, we set $\mathtt{accumulate\_grad\_batches}$ to 16 to achieve an effective batch size of 16.
Training loop Each epoch of the training loop involves training the model with a maximum of 400 batches from the training set, followed by computing the NLL on the validation set. Training stops when any of these conditions are met:
  • We’ve reached a total of 10,000 gradient updates during model training,
  • We observed no improvement in the best value of the NLL on the validation set for 10 consecutive epochs.
We return the version of the model that achieved the best value of the NLL on the validation set.
@@ A.4 Naive Baseline Description
In this paper, we employ VAR [33] (Vector Autoregression) and GARCH [34] (Generalized Autoregressive Conditionally Heteroscedastic) as two naive baseline models. The VAR($p$) model is defined as:
@@
where $A_{i}$ is a $N\times N$ coefficient matrix and $\mathbf{c}$ is the intercept. We use a VAR model of lag 1 (i.e., a VAR(1) model) in the experiments. The parameters of Eq. (19) are estimated using ordinary least squares (OLS) following the procedure in [33].
The GARCH model describes the conditional covariance matrix of the current error term of a multivariate system. Suppose the model for the conditional mean is an AR($1$) model:
@@
where the error term is modeled as:
@@
where $\boldsymbol{H}_{t}$ is a $N\times N$ conditional covariance matrix and $\mathbf{e}_{t}$ is a $N\times 1$ standard normal vector $\mathbf{e}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$. In the experiments, we use the DCC-GARCH(1,1)11(1,1)( 1 , 1 ) model [35], where the conditional covariance matrix $\boldsymbol{H}_{t}$ is defined as:
@@
where $\boldsymbol{D}_{t}=\operatorname{diag}{(\mathbf{h}_{t})}^{1/2}$, and $\mathbf{h}_{t}$ contains the variances for each time series. $\boldsymbol{R}_{t}$ is the conditional correlation matrix in DCC-GARCH. The parameters of the DCC-GARCH model are estimated with the log-likelihood function:
@@
In this paper, we implement the VAR model with $\mathtt{statsmodels}$ [36] and the DCC-GARCH model with $\mathtt{mgarch}$ [37].
@@ A.5 Hyperparameter Search
The hyperparameters and training configuration largely align with those in the GPVar paper [2]. The parameters of all DL models are learned using the Adam optimizer with $l2$ regularization set to 1e-8, and gradients are clipped to 10.0. Across all methods, we limit the total number of gradient updates to 10,000, and decay the learning rate by a factor of 2 after 500 consecutive updates without improvement. Table 3 lists the parameters that are tuned, along with the values of hyperparameters that are kept constant across all datasets and not subject to tuning.
@@
To tune the hyperparameters of each model, we conduct a grid search of 9 parameters on each dataset. The best hyperparameters for a base model—dataset combination are selected based on the best validation loss. Once the optimal learning rate and hidden size are determined, we apply the same hyperparameters for models both with and without applying our approach.
The number of base kernels ($M$) for generating $\boldsymbol{C}_{t}$ and the associated lengthscale set $\{l_{m}\}_{m=1}^{M-1}$ are two additional hyperparameters when applying our method. The optimal values of $M$ and $\{l_{m}\}_{m=1}^{M-1}$ are selected in a similar way with hyperparameter search. The values of $M$ and $\{l_{m}\}_{m=1}^{M-1}$ searched during hyperparameter tuning are shown in Table 4. There are 6 possible combinations. For example, if we set $M=3$ and choose the lengthscale to start at $1.0$, the lengthscales for generating the component kernels will be $\{1.0,2.0\}$ since the last weight is for the identity matrix.
@@ A.6 Base Model Description and Input Features
The input to the base models consists of lagged time series values and generic features that encode time and time series identification. The number of lagged values used is determined by the time-frequency of each dataset. Specifically, we use [1, 24, 168] for hourly data, [1, 7, 14] for daily data, and [1, 2, 4, 12, 24, 48] for data with a granularity of less than one hour. For all other datasets, we only use the lag-1 values.
We use generic features to represent time. For datasets with a granularity of one hour or less, we include hour of the day and day of the week features. For daily datasets, we use the day of the week feature. Additionally, each time series is distinguished by an identifier number. All features are encoded with a single value; for example, the hour of the day feature takes values in [0, 23]. These feature values are concatenated with the RNN or Transformer input at each time step to generate the model input vector $\mathbf{y}_{t}$.
As illustrated in §4, our method requires a state vector $\mathbf{h}_{t}$ at each time step to generate the parameters for the predictive distribution and the dynamic weights for correlation matrix kernels. We use two different neural architectures for this purpose: RNN and Transformer, both of which preserve autoregressive properties. Specifically, we use LSTM as our base model for RNN and a decoder-only Transformer (i.e., the GPT model [22]) for the Transformer.
LSTM, a type of RNN architecture, is designed to model sequences and time series data. Unlike traditional RNNs, LSTMs can learn long-term dependencies, making them effective for tasks requiring context and memory over long sequences. A decoder-only Transformer is primarily used for sequence generation tasks, such as text generation, language modeling, and machine translation. It is a simplified version of the original Transformer model introduced by Vaswani et al. [38], including only the decoder part. The LSTM model can be formulated as:
@@
where $\mathbf{y}_{t}$ is the input at each time step. The decoder-only Transformer can be formulated as:
@@ B.1 Metric Definition
In this paper, we repeat the evaluation process on the testing set 10 times to compute the mean and standard deviation of all metrics. Metrics calculated in each independent evaluation process are based on the average results from all forecast instances in the testing set. For example, the $\operatorname{CRPS}_{\text{sum}}$ reported for $\mathtt{traffic}$ will be the average $\operatorname{CRPS}_{\text{sum}}$ of 7 forecast instances in its testing set. 100 samples of prediction were drawn for all evaluation processes.
@@ B.1.1 Continuous Ranked Probability Score
where we use the empirical CDF to represent $F_{i,t}$ when predicting $z_{i,t}$. Since $\operatorname{CRPS}$ only compares a single ground-truth value to its predicted distribution, we also calculate the $\operatorname{CRPS}_{\text{sum}}$ [2, 16, 16] to assess multivariate dependencies in the time series data. $\operatorname{CRPS}_{\text{sum}}$ is computed by summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ over the resulting sums:
@@ B.1.2 Quantile Loss
where $\mathrm{I}$ is a binary indicator function that equals 1 when the condition is met, $\hat{z}^{\rho}$ represents the predicted $\rho$-quantile, and $z$ represents the ground truth value. The quantile loss serves as a metric to assess the accuracy of a given quantile, denoted by $\rho$, from the predictive distribution. We summarize the quantile losses for the testing set across all time series segments by evaluating a normalized summation of these losses: $\left(\sum_{i,t}L_{\rho}\left(z_{i,t},\hat{z}_{i,t}^{\rho}\right)\right)/\left(\sum_{i,t}z_{i,t}\right)$. In this paper, we evaluate the $0.5$-risk and the $0.9$-risk following Salinas et al. [9].
@@ B.1.3 Energy Score
The energy score ($\operatorname{ES}$) generalizes the $\operatorname{CRPS}$ to evaluate distributional forecasts of a vector-valued random variate, and thus another multivariate metric used in this paper:
@@
where $\lVert\mathbf{z}\rVert_{2}$ is the Euclidean norm and we use $\beta=1$ following [16] in this paper. Since we also want to aggregate over the prediction horizon, we calculate the Frobenius norm of the matrix $\lVert\mathbf{z}_{t+1:t+Q}\rVert_{F}$ in practice.
@@ B.1.4 Root Relative Mean Squared Error
The root relative mean squared error ($\operatorname{RRMSE}$) is a metric commonly used for point forecast [39, 27, 40]. $\operatorname{RRMSE}$ is defined as:
@@
where $\hat{\mathbf{z}}_{t}$ is obtained by taking the mean of our prediction samples, $\bar{\mathbf{z}}$ is the mean value of the whole forecast instance. We use this metric to evaluate the mean prediction performance of our model.
@@ B.2 Results on Other Forecasting Metrics
The results for $\operatorname{CRPS}$ (Table 5), $0.5$-risk (Table 6), $0.9$-risk (Table 7), $\operatorname{ES}$ (Table 8), and $\operatorname{RRMSE}$ (Table 9) are presented. An "N/A" entry in the tables indicates that the naive baseline models could not fit properly for this dataset. We observe consistent performance improvements in the base models using our method across different evaluation metrics. Notably, in the multivariate metric $\operatorname{ES}$, our method shows significant improvement, with an average increase of 5.58% for GPVar and 3.21% for Transformer.
@@ B.3 Training Dynamics
In Fig. 4 and Fig. 5, we compare the training dynamics of the base models trained with and without our method. Note that the likelihood losses of the two methods are not directly comparable, even for the same dataset and base model, due to differences in likelihood structure. We observe that while our method introduces more complexity into the likelihood loss, there is no evidence that it prolongs model convergence. On the contrary, our method can speed up convergence for some datasets in terms of the training steps used. We also report the training time in Table 10.
@@
B.4 Additional Model Interpretation
In this section, we provide further insights into how our method improves the base model. We illustrate these insights by comparing the cross-correlation of the residuals from models with and without our method. Additionally, we demonstrate the performance of our method over the prediction horizon in multistep-ahead forecasting.
B.4.1 Comparison of Residual Correlation
Recall that our method models both the autocovariance of errors $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{i,t})$ and the cross-lag covariance $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{j,t})$ between all pairs of components in the multivariate series. With the calibration process introduced in §4.2, our method is expected to reduce error cross-correlation (autocorrelation and cross-lag correlation). Here, we compare the empirical ACF of the residuals $\eta_{i,t}$ of a single time series $i$, as well as the empirical cross-correlations of $\boldsymbol{\eta}_{t}$ across multiple time series.
@@ B.4.1 Comparison of Residual Correlation
Next we compare cross-correlations of the one-step-ahead prediction residuals with and without our method. The comparisons are provided for the following datasets: $\mathtt{electricity}$ (Fig. 13), $\mathtt{traffic}$ (Fig. 14), $\mathtt{wiki}$ (Fig. 15), $\mathtt{m4\_hourly}$ (Fig. 16), $\mathtt{pems03}$ (Fig. 17), and $\mathtt{uber\_hourly}$ (Fig. 18). We also observe that the cross-correlations of the residuals are reduced after applying our method.
@@
B.4.2 Performance Breakdown at Each Forecast Step
To investigate our performance gain at each forecast step, we calculate the $\operatorname{CRPS}_{\text{sum}}$ for each forecast step. The results are shown in Fig. 19 (GPVar) and Fig. 20 (Transformer). Note that the $\operatorname{CRPS}_{\text{sum}}$ reported in this section may have different scales compared to previous sections since they are not normalized. Since the predicted values are used as inputs within the prediction range during multistep-ahead forecasting, the residuals accumulate the effects of inaccuracies from previous steps. Therefore, the performance improvement depends not only on our modeling of error correlations but also on the properties of the residuals. These properties can be influenced by the absolute and relative time of the forecast and the seasonality of the data. For data without strong seasonality, residuals tend to be larger when predicting further ahead, making error accumulation more apparent. In contrast, for data with strong seasonality, the impact can vary. We can observe that, in most scenarios, $\operatorname{CRPS}_{\text{sum}}$ is reduced at the early forecasting stages. As predictions extend further into the future, some datasets (e.g., $\mathtt{traffic}$ in Fig. 19) show decreased improvement, likely due to seasonality effects. In contrast, other datasets (e.g., $\mathtt{wiki}$ in Fig. 19) exhibit larger improvements further into the future, possibly because the residuals accumulate over the steps.

### v4

@@ A.3 Training Procedure
Compute used All models in the paper were trained in an Anaconda environment with access to one AMD Ryzen Threadripper PRO 5955WX CPU and four NVIDIA RTX A5000 GPUs (each with 24 GB of memory).
Batch size We adopt the approach of GPVar [3] by using $B=20$ time series in a sample slice and a batch size of 16. Because our data sampler selects one slice of time series as a batch instead of sampling 16 slices simultaneously, we set $\mathtt{accumulate\_grad\_batches}$ to 16 to achieve an effective batch size of 16.
Training loop Each epoch involves training the model on up to 400 batches from the training set, followed by computing the NLL on the validation set. Training stops when any of the following conditions are met:
  • A total of 10,000 gradient updates have been performed during model training,
  • No improvement in the best NLL value on the validation set is observed for 10 consecutive epochs.
We select the version of the model that achieved the best NLL value on the validation set.
@@ A.4 Naive Baseline Description
In this paper, we employ VAR [38] (Vector Autoregression) and GARCH [39] (Generalized Autoregressive Conditionally Heteroskedasticity) as two naive baseline models. The VAR($p$) model is defined as
@@
where $A_{i}$ is an $N\times N$ coefficient matrix, and $\mathbf{c}$ is the intercept. We use a VAR model of lag 1 (i.e., a VAR(1) model) in the experiments. The parameters of Eq. (19) are estimated using ordinary least squares (OLS), following the procedure in [38].
The GARCH model describes the conditional covariance matrix of the error term in a multivariate system. Suppose the model for the conditional mean is an AR($1$) model:
@@
where the error term is modeled as
@@
where $\boldsymbol{H}_{t}$ is an $N\times N$ conditional covariance matrix, and $\mathbf{e}_{t}$ is an $N\times 1$ standard normal vector, $\mathbf{e}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$. In the experiments, we use the DCC-GARCH(1,1)11(1,1)( 1 , 1 ) model [40], where the conditional covariance matrix $\boldsymbol{H}_{t}$ is defined as
@@
where $\boldsymbol{D}_{t}=\operatorname{diag}{(\mathbf{h}_{t})}^{1/2}$, and $\mathbf{h}_{t}$ contains the variances for each time series. $\boldsymbol{R}_{t}$ is the conditional correlation matrix in the DCC-GARCH model. The parameters of the DCC-GARCH model are estimated with the log-likelihood function:
@@
In this paper, we implement the VAR model using $\mathtt{statsmodels}$ [41] and the DCC-GARCH model using $\mathtt{mgarch}$ [42].
@@ A.5 Hyperparameter Search
The hyperparameters and training configuration largely align with those used in the GPVar paper [3]. All DL models are trained using the Adam optimizer with $l2$ regularization set to 1e-8, and gradients are clipped at 10.0. For all methods, we limit the total number of gradient updates to 10,000 and decay the learning rate by a factor of 2 after 500 consecutive updates without improvement. Table 3 lists the parameters that are tuned, as well as the hyperparameters that are kept constant across all datasets and not subject to tuning.
@@
To tune the hyperparameters of each model, we conduct a grid search over nine parameters on each dataset. The best hyperparameters for each base model–dataset combination are selected based on the lowest validation loss. Once the optimal learning rate and hidden size are determined, we apply the same hyperparameters to models both with and without our method.
The number of base kernels ($M$) for generating $\boldsymbol{C}_{t}$ and the associated lengthscale set $\{l_{m}\}_{m=1}^{M-1}$ are two additional hyperparameters when applying our method. The optimal values of $M$ and $\{l_{m}\}_{m=1}^{M-1}$ are selected in a similar manner via hyperparameter search. The values of $M$ and $\{l_{m}\}_{m=1}^{M-1}$ explored during hyperparameter tuning are shown in Table 4. There are six possible combinations. For example, if we set $M=3$ and choose the initial lengthscale to be $1.0$, the lengthscales for generating the component kernels will be $\{1.0,2.0\}$ since the last weight corresponds to the identity matrix.
@@ A.6 Base Model Description and Input Features
The input to the base models consists of lagged time series values and generic features that encode time and identify each time series. The number of lagged values used is determined by the time-frequency of each dataset. Specifically, we use lags [1, 24, 168] for hourly data; [1, 7, 14] for daily data; and [1, 2, 4, 12, 24, 48] for data with a granularity of less than one hour. For all other datasets, we only use the lag-1 values.
We use generic features to represent time. For datasets with a granularity of one hour or less, we include features for the hour of the day and the day of the week. For daily datasets, we use the day of the week feature. Additionally, each time series is distinguished by an identifier number. All features are encoded with a single value; for example, the hour of the day feature takes values in [0, 23]. These feature values are concatenated with the RNN or Transformer input at each time step to generate the model input vector $\mathbf{y}_{t}$.
As illustrated in §4, our method requires a state vector $\mathbf{h}_{t}$ at each time step to generate the parameters for the predictive distribution and the dynamic weights for correlation matrix kernels. We use two different neural architectures for this purpose: RNN and Transformer, both of which preserve autoregressive properties. Specifically, we use an LSTM as our base model for the RNN and a decoder-only Transformer (i.e., the GPT model [25]) for the Transformer. Table 5 and Table 6 summarize the number of parameters for the GPVar and Transformer models across each dataset.
Table 5: Number of parameters of the GPVar model for each dataset.
covariate embedding rnn distribution proj covariance proj (our method)
$\mathtt{exchange\_rate}$ 60 6.1k 252 84
$\mathtt{solar}$ 3.7k 26.6k 492 164
$\mathtt{electricity}$ 16.5k 29.6k 492 164
$\mathtt{traffic}$ 72.5k 34.6k 492 164
$\mathtt{wikipedia}$ 200k 5.7k 132 44
$\mathtt{m4\_hourly}$ 19.7k 10.2k 252 84
$\mathtt{m1\_quarterly}$ 6.3k 25k 492 164
$\mathtt{pems03}$ 26.4k 34.6k 492 164
$\mathtt{uber\_hourly}$ 9.7k 28.3k 492 164
Table 6: Number of parameters of the Transformer model for each dataset.
target proj covariate proj covariate embedding transformer distribution proj covariance proj (our method)
$\mathtt{exchange\_rate}$ 160 400 60 26.5k 492 164
$\mathtt{solar}$ 40 400 3.7k 1.8k 132 44
$\mathtt{electricity}$ 160 2.4k 16.5k 26.5k 492 164
$\mathtt{traffic}$ 80 1.8k 72.5k 6.8k 252 84
$\mathtt{wikipedia}$ 160 4.2k 200k 26.5k 492 164
$\mathtt{m4\_hourly}$ 80 1.2k 19.7k 6.8k 252 84
$\mathtt{m1\_quarterly}$ 80 1.3k 6.3k 26.5k 492 164
$\mathtt{pems03}$ 70 870 26.4k 1.8k 132 44
$\mathtt{uber\_hourly}$ 160 2k 9.7k 26.5k 492 164
LSTM, a type of RNN architecture, is designed to model sequences and time series data. Unlike traditional RNNs, LSTMs can learn long-term dependencies, making them effective for tasks requiring context and memory over long sequences. A decoder-only Transformer is primarily used for sequence generation tasks, such as text generation, language modeling, and machine translation. It is a simplified version of the original Transformer model introduced by Vaswani et al. [43], consisting of only the decoder component. The LSTM model can be formulated as
@@
where $\mathbf{y}_{t}$ is the input at each time step. The decoder-only Transformer can be formulated as
@@ B.1 Metric Definition
In this paper, we repeated the evaluation process on the testing set ten times to compute the mean and standard deviation of all metrics. Metrics calculated in each independent evaluation are based on the average results from all forecast instances in the testing set. For example, the $\operatorname{CRPS}_{\text{sum}}$ reported for $\mathtt{traffic}$ is the average $\operatorname{CRPS}_{\text{sum}}$ of seven forecast instances in its testing set. 100 prediction samples were drawn for all evaluation processes.
@@ B.1.1 Continuous Ranked Probability Score
where we use the empirical CDF to represent $F_{i,t}$ when predicting $z_{i,t}$. Since $\operatorname{CRPS}$ only compares a single ground-truth value to its predicted distribution, we also calculate the $\operatorname{CRPS}_{\text{sum}}$ [3, 16, 17] to assess multivariate dependencies in the time series data. $\operatorname{CRPS}_{\text{sum}}$ is computed by summing both the forecasted and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ over the resulting sums:
@@ B.1.2 Quantile Loss
where $\mathrm{I}$ is a binary indicator function that equals 1 when the condition is met, $\hat{z}^{\rho}$ represents the predicted $\rho$-quantile, and $z$ represents the ground truth value. The quantile loss serves as a metric to assess the accuracy of a given quantile $\rho$ from the predictive distribution. We summarize the quantile losses over the testing set across all time series segments by computing a normalized summation of these losses: $\left(\sum_{i,t}L_{\rho}\left(z_{i,t},\hat{z}_{i,t}^{\rho}\right)\right)/\left(\sum_{i,t}z_{i,t}\right)$. In this paper, we evaluate the $0.5$-risk and the $0.9$-risk following Salinas et al. [10].
@@ B.1.3 Energy Score
The Energy Score ($\operatorname{ES}$) generalizes the $\operatorname{CRPS}$ to evaluate distributional forecasts of a vector-valued random variable and is thus another multivariate metric used in this paper:
@@
where $\lVert\mathbf{z}\rVert_{2}$ is the Euclidean norm. In this paper, we use $\beta=1$, following [17]. Since we also want to aggregate over the prediction horizon, we calculate the Frobenius norm of the matrix $\lVert\mathbf{z}_{t+1:t+Q}\rVert_{F}$ in practice.
@@ B.1.4 Root Relative Mean Squared Error
The Root Relative Mean Squared Error ($\operatorname{RRMSE}$) is a metric commonly used for point forecasts [44, 32, 45]. $\operatorname{RRMSE}$ is defined as:
@@
where $\hat{\mathbf{z}}_{t}$ is obtained by taking the mean of our prediction samples, and $\bar{\mathbf{z}}$ is the mean value of the entire forecast instance. We use this metric to evaluate the mean prediction performance of our model.
@@ B.2 Results on Other Forecasting Metrics
We present the results for $\operatorname{CRPS}$ (Table 7), the $0.5$-risk (Table 8), the $0.9$-risk (Table 9), $\operatorname{ES}$ (Table 10), and $\operatorname{RRMSE}$ (Table 11). An “N/A” entry in the tables indicates that the naive baseline models could not be properly fitted to this dataset. We observe consistent performance improvements in the base models using our method across different evaluation metrics. Notably, in the multivariate metric $\operatorname{ES}$, our method shows significant improvement, reducing the score by an average of 5.58% for GPVar and 3.21% for the Transformer.
@@ B.3 Training Dynamics
In Fig. 5 and Fig. 6, we compare the training dynamics of the base models trained with and without our method. Note that the likelihood losses of the two methods are not directly comparable, even for the same dataset and base model, due to differences in the likelihood structures. We observe that while our method introduces more complexity into the likelihood function, there is no evidence that it significantly prolongs model convergence. On the contrary, our method can speed up convergence for some datasets in terms of the training steps used. We also report the training time in Table 12.
@@
B.4 Effect of the Number of Time Series during Prediction
The number of time series does not impact training, as the model is trained using a random subset of $B$ time series at a time, independent of the total number of time series $N$. However, during prediction, the batch size can be increased beyond the training batch size of $B=20$ for multistep-ahead rolling predictions. This allows more information to be utilized, potentially improving both predictions and error calibration, provided that memory capacity permits. We conducted an additional experiment to demonstrate the effect of increasing the batch size during inference (Fig. 7).
Figure 7: The influence of the number of time series in a batch on the performance of inference. "w/o" denotes methods without time-dependent errors, while "w/" indicates our method. We only show some datasets here because the remaining datasets have fewer than $B=20$ time series in the testing set.
B.5 Additional Model Interpretation
In this section, we provide further insights into how our method improves the base model. We illustrate these improvements by comparing the cross-correlations of the residuals from models with and without our method. Additionally, we demonstrate the performance of our method over the prediction horizon in multistep-ahead forecasting.
B.5.1 Comparison of Residual Correlation
Recall that our method models both the autocovariance of errors $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{i,t})$ and the cross-lag covariance $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{j,t})$ between all pairs of components in the multivariate series. With the calibration process introduced in §4.2, our method is expected to reduce error cross-correlations, including autocorrelation and cross-lag correlation. Here, we compare the empirical ACF of the residuals $\eta_{i,t}$ of a single time series $i$, as well as the empirical cross-correlations of $\boldsymbol{\eta}_{t}$ across multiple time series.
@@ B.5.1 Comparison of Residual Correlation
Next, we compare the cross-correlations of the one-step-ahead prediction residuals with and without our method. The comparisons are provided for the following datasets: $\mathtt{electricity}$ (Fig. 15), $\mathtt{traffic}$ (Fig. 16), $\mathtt{wiki}$ (Fig. 17), $\mathtt{m4\_hourly}$ (Fig. 18), $\mathtt{pems03}$ (Fig. 19), and $\mathtt{uber\_hourly}$ (Fig. 20). We also observe that the cross-correlations of the residuals are reduced after applying our method.
@@
B.5.2 Performance Breakdown at Each Forecast Step
To investigate our performance gain at each forecast step, we calculate the $\operatorname{CRPS}_{\text{sum}}$ for each forecast step. The results are shown in Fig. 21 for GPVar and Fig. 22 for the Transformer. Note that the $\operatorname{CRPS}_{\text{sum}}$ reported in this section may have different scales compared to previous sections because they are not normalized. In multistep-ahead forecasting, since the predicted values are used as inputs for subsequent predictions within the prediction range, the residuals accumulate the effects of inaccuracies from previous steps. Therefore, the performance improvement depends not only on our modeling of error correlations but also on the properties of the residuals. These properties can be influenced by the absolute and relative time of the forecast and the seasonality of the data. For data without strong seasonality, residuals tend to be larger when predicting further ahead, making error accumulation more apparent. Conversely, for data with strong seasonality, the impact of error accumulation can vary. We observe that, in most scenarios, $\operatorname{CRPS}_{\text{sum}}$ is reduced at the early forecasting stages. As predictions extend further into the future, some datasets (e.g., $\mathtt{traffic}$ in Fig. 21) show decreased improvement, likely due to seasonality effects. Conversely, other datasets (e.g., $\mathtt{wiki}$ in Fig. 21) exhibit larger improvements further into the future, possibly because the residuals accumulate over the steps.